For the test of significance questions, clearly indicate each of the formal steps in the test of significance.

Step 1: State the null and alternative hypothesis.
Step 2: Calculate the test statistic.
Step 3: Find the p-value.

Step 4: State your conclusion. (Do not just say “Reject H0” or “Do not reject H0”, state the conclusion in the context of the problem.)

In a recent survey of county high school students, 100 males and 100 females, 66 of the male students and 47 of the female students sampled admitted that they consumed alcohol on a regular basis. Find a 90% confidence interval for the difference between the proportion of male and female students that consume alcohol on a regular basis. Can you draw any conclusions from the confidence interval?
Answer #1

Here we have n=100 m 100 22 247 24 66 < X2 m 47 100 66 -B Els 100 0.47 20.66 0.565 Ia + 2 66 +47 100 +100 n, + m Standard Err

Steph: Hypothesis Ho: P - P2 = 0 Ha: P - P2 & 0 This is a two failed test Stepair Pext statistics - 0.66 -0.47 S.A 0.070) Za

20.66-047 = 1,645 * 0.0701 At 90, confidence intervals , ♡ - 1-0.90 = 0.10 a12 20.05 Zcritical a 20.05 2 1.6455. 9ou confide

There is sufficient evidence to supplert the claim that there is difference between the Proportion of male and females studen

Dear students , please like it.

Thanks.